OpenStudy (anonymous):

Simplify 2x + 10 --------- 2 A. x B. x + 10 C. 2x + 5 D. x+ 5

OpenStudy (lgbasallote):

\[\large \frac{2x}{2} + \frac{10}{2}\] you can divide eacch term of the numerator by the denominator

OpenStudy (anonymous):

I still don't get...

OpenStudy (anonymous):

\[\frac{2x+10}{2}=\frac{2x}{2}+ \frac{10}{2}\ = x+5 \]

OpenStudy (anonymous):

Thanks, Luis.

OpenStudy (anonymous):

kk